# On anomalous dissipation induced by transport noise

### v2

@@ Abstract.
In this paper, we show that suitable transport noises produce anomalous dissipation of enstrophy and energy of solutions of the 2D Navier-Stokes equations and diffusion equations in all dimensions, respectively. The key ingredients are Meyers’ type estimates for SPDEs with transport noise which are combined with recent scaling limits for such SPDEs. The former allows us to obtain, for the first time for such type of scaling limits, convergence in a space of positive smoothness uniformly in time. Compared to known results, one of the main novelties is that anomalous dissipation might take place even in the presence of a transport noise of arbitrarily small intensity. Further, we discuss physical interpretations.
@@ 1. Introduction and statement of the main results
(1.1) $\partial_{t}\zeta^{\nu}&+(u^{\nu}\cdot\nabla)\zeta^{\nu}=\nu\Delta\zeta^{\nu}+\sqrt{2\mu}\sum_{k\in{\mathbb{Z}}^{2}_{0}}\theta_{k}^{\nu}\,(\sigma_{k}\cdot\nabla)\zeta^{\nu}\circ\dot{W}_{t}^{k}&\text{on }&\mathbb{T}^{2},\\ \zeta^{\nu}(0,\cdot)&=\zeta_{0}&\text{on }&\mathbb{T}^{2},\end{aligned}\right.$
where $u^{\nu}={\mathcal{K}}\,\zeta^{\nu}\stackrel{{\scriptstyle{\rm def}}}{{=}}(-\partial_{y},\partial_{x})\Delta^{-1}\zeta^{\nu}$, ${\mathcal{K}}\,$ denotes the Biot-Savart operator, $\mathbb{T}^{2}$ the two-dimensional torus, $\nu>0$ the kinematic viscosity of the fluid, $(W^{k})_{k\in{\mathbb{Z}}^{2}_{0}}$ a family of complex Brownian motions, $\circ$ the Stratonovich product, $\theta^{\nu}=(\theta_{k}^{\nu})_{k\in{\mathbb{Z}}^{2}_{0}}\in\ell^{2}$ with ${\mathbb{Z}}^{2}_{0}={\mathbb{Z}}^{2}\setminus\{0\}$, $(\sigma_{k})_{k\in{\mathbb{Z}}^{2}_{0}}$ a family of divergence-free vector field described in Subsection 2.2, and $\mu>0$ the noises intensity (cf., (1.5) below).
Anomalous energy dissipation in fluid flows has been proven experimentally to a large degree [78] and stands at the basis of turbulence theory [43, 53, 54, 55]. For this reason, it is sometimes referred to as the zeroth law of turbulence. Roughly speaking, anomalous dissipation states that at high Reynolds numbers (i.e., $\nu\downarrow 0$) the averages of the energy dissipation are uniformly bounded from below in $\nu$. In the context of the Kraichnan-Batchelor theory (KB in the following) of 2D turbulence [13, 56] (see also [43, Section 9.7]) the relevant anomalously dissipated quantity is the enstrophy and the anomalous dissipation of enstrophy reads as:
@@
[…]
In the deterministic setting (i.e., $\theta^{\nu}\equiv 0$), there are serious obstructions in making the above pictures rigorous. Indeed, as explained in [60, Section 1], in the case of $\zeta_{0}\in L^{2}(\mathbb{T}^{2})$, the anomalous dissipation of enstrophy (1.2) implies the existence of enstrophy-dissipative (suitable) weak solutions to the deterministic incompressible 2D Euler equations obtained via vanishing viscosity limits. However, the latter is known not to be true, see e.g., [27, 70] or [25, Remark 2]. Moreover, the proposed ways to fix such paradox in [31] appeared to not be resolutive [60, Section 7].
The anomalous dissipation result for the enstrophy for the 2D NSEs (1.1) in Theorem 1.1 below appears to be the first result in the direction of the KB theory in the physically relevant case of $L^{2}(\mathbb{T}^{2})$ initial enstrophy, see also the comments below (1.7). However, here we are not aiming at capturing the sophisticated mechanics behind anomalous dissipation as described below (1.2). Instead, we prove that transport noises, localised at high frequencies and varying with the viscosity, lead to anomalous dissipation. More precisely, anomalous dissipation is created by the ‘turbulent flows’ $\sum_{k\in{\mathbb{Z}}^{2}_{0}}\theta^{\nu}_{k}\sigma_{k}\,\dot{W}^{k}$ rather than the convective nonlinearity as expected in the KB theory. Moreover, this is also true for noises of ‘small intensities’ and for ‘nonlinear passive scalar’ such as the 2D NSEs in vorticity formulation (here, we mean the turbulent flows differ from the advected quantity). Surprisingly, stochastic perturbations of 2D NSEs provide new dissipation mechanics compared to the deterministic case. The reader is also referred to Section 5 where we also discuss the velocity formulation of the NSEs and three-dimensional case. Further details about physical interpretations of our results are given in Subsection 1.2.
@@
(1.3) $\partial_{t}\varrho^{\gamma}&=\gamma\Delta\varrho^{\gamma}+\sqrt{c_{d}\mu}\sum_{k\in{\mathbb{Z}}^{d}_{0}}\sum_{1\leq\alpha\leq d-1}\theta_{k}^{\gamma}\,(\sigma_{k,\alpha}\cdot\nabla)\varrho^{\gamma}\circ\dot{W}_{t}^{k,\alpha}&\text{on }&\mathbb{T}^{d},\\ \varrho^{\gamma}(0,\cdot)&=\varrho_{0}&\text{on }&\mathbb{T}^{d}.\end{aligned}\right.$
@@
Here, $\mathbb{T}^{d}$ is the $d$-dimensional torus, $\gamma>0$ the diffusivity of the scalar, $\theta^{\nu}=(\theta_{k}^{\gamma})_{k\in{\mathbb{Z}}^{d}_{0}}\in\ell^{2}$, $c_{d}=\frac{d}{d-1}$, while $(W^{k,\alpha})_{k,\alpha}$ and $(\sigma_{k,\alpha})_{k,\alpha}$ are families of complex Brownian motions and divergence-free vector fields described in Subsection 2.2.
@@
Anomalous dissipation for passive scalars advected by turbulent flows has attracted a lot of interest in recent years. For passive scalars, anomalous dissipation by turbulent flows is at the basis of the corresponding theory of scalar turbulence [29, 76, 77]. It can be defined by replacing in (1.2) the viscosity $\nu$ and the vorticity $\zeta^{\nu}$ by the diffusivity $\gamma>0$ and the intensity of the passive scalar $\varrho^{\gamma}$ (solving an advection-diffusion such as (1.3)), respectively. In the deterministic setting, many results have been established in the context of anomalous dissipation. It is not possible to give here a complete overview of the deterministic results and the reader is referred to, e.g., [12, 17, 21, 24, 26, 30] and the references therein.
@@
Before discussing works on anomalous dissipation and related topics in the context of stochastic fluid dynamics, let us first discuss the physical relevance of transport noise in stochastic fluid dynamics. Nowadays there are several derivations of NSEs with transport noise available in the literature, see e.g., [28, 41, 42, 50], and it is by now a well-established model in stochastic fluid dynamics [18, 19, 20, 23, 35, 40, 47, 48]. To some extent, at the basis of such derivations is the idea of separation of scales. A heuristic derivation using the latter principle is given in Subsection 1.2 below. This allows us to stress the physical relevance of the $\nu$-dependence of transport noise in (1.1) and provide an interpretation of our results for 2D NSEs. Starting from the works [37, 44], the effect of transport noise on mixing and enhanced dissipation of NSEs or advection-diffusion equations is by now well-understood, see e.g., [34, 61, 33] (see also [1, 22, 32, 36, 38, 39, 58, 62] for related works). The same is not true for anomalous dissipation. To the best of our knowledge, the only results on the anomalous dissipation for SPDEs with transport noise are given in [49, 73]. A comparison is postponed to Subsection 1.3.4 where a more detailed discussion on our contribution is possible. Finally, from a technical point of view, our proofs rely on a refinement of the scaling limit arguments [37, 44] in the parabolic setting which is interesting on its own. A detailed discussion is given in Subsection 1.3.2 below.
@@ 1.1. Anomalous dissipation results
Our main results show that, with a suitable transport noise, anomalous dissipation holds with:
  • small noise – Theorems 1.1(1) and 1.2(1);
  • prescribed rate – Theorems 1.1(2) and 1.2(2).
As highlighted above, compared to the existing literature, the main contribution of the current work is to reveal that anomalous dissipation can take place even in the presence of a small transport noise; while previous works typically require noise with high intensity (as above, details are given in Subsection 1.3.4).
In the statement of our main results, we employ the following notation
@@
Let $N\geq 1$ and $\delta>0$ be fixed. Then the following hold.
  1. (1)
    (Small noise). For all $\mu>0$, there exists a family $(\theta^{\nu})_{\nu\in(0,1)}\subseteq\mathcal{S}_{\ell^{2}}^{0}$ such that, for all mean-zero $\zeta_{0}\in H^{\delta}(\mathbb{T}^{2})$ satisfying $N^{-1}\leq\|\zeta_{0}\|_{L^{2}}$ and $\|\zeta_{0}\|_{H^{\delta}}\leq N$, we have
    $\inf_{\nu\in(0,1)}\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{2}}\nu\,|\nabla\zeta^{\nu}|^{2}\,\mathrm{d}x\,\mathrm{d}t>0.$
  2. (2)
    (Prescribed rate). For all $\eta\in(0,1)$ and $\mu>-\ln(1-\eta)$, there exists a family $(\theta^{\nu})_{\nu\in(0,1)}\subseteq\mathcal{S}_{\ell^{2}}^{0}$ such that, for all mean-zero $\zeta_{0}\in H^{\delta}(\mathbb{T}^{2})$ satisfying $N^{-1}\leq\|\zeta_{0}\|_{L^{2}}$ and $\|\zeta_{0}\|_{H^{\delta}}\leq N$, we have
    $\inf_{\nu\in(0,1)}\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{2}}\nu\,|\nabla\zeta^{\nu}|^{2}\,\mathrm{d}x\,\mathrm{d}t\geq\frac{\eta}{2}\,\|\zeta_{0}\|_{L^{2}}^{2}.$
In the statements in (1)-(2), $\zeta^{\nu}$ denotes the unique global smooth solution to (1.1).
@@
Since Theorem 1.1(1) holds for all $\mu>0$, it indeed corcerns ‘small’ noises. Instead, Theorem 1.1(2) ensures that one can prescribe the rate of dissipation at the expense of considering a uniformly $\nu$-bounded with large intensity transport noise. The case $\zeta_{0}=0$ is excluded as no anomalous dissipation can take place. Existence of global smooth solutions of (1.1) with $\theta^{\nu}\in\mathcal{S}_{\ell^{2}}^{0}$ follows as in [9, Theorems 2.7 and 2.12].
An inspection of the proof of Theorem 1.1(1)-(2) shows that the time $t=1$ can be replaced by any time $t>0$; however, the corresponding choice of $(\theta^{\nu})_{\nu\in(0,1)}$ depends on such time $t$. In addition, we can take $\theta^{\nu}$ to be constants for $\nu\in[\nu_{j+1},\nu_{j})$ where $(\nu_{j})_{j\geq 1}\subseteq(0,1]$ is a sequence satisfying $\nu_{1}=1$, $\nu_{j+1}\leq\nu_{j}$ and $\lim_{j\to\infty}\nu_{j}=0$. Finally, the proof of Theorem 1.1 shows that $\theta^{\nu}$’s are localized at high frequencies:
@@
To connect Theorem 1.1 with the discussion around (1.2) on KB theory, let us discuss its implication. Firstly, as $\nabla\cdot\sigma_{k,\alpha}=0$, solutions to (1.1) satisfy:
If $\theta^{\nu}$ is chosen as in Theorem 1.1, then the 2D NSEs (1.1) exhibit anomalous dissipation of enstrophy at time $t=1$, i.e.,
Thus, with the aid of transport noise, we obtain anomalous dissipation of enstrophy while keeping true the energy balance (1.7) at fixed $\nu>0$ and with $\zeta_{0}\in L^{2}(\mathbb{T}^{2})$. As mentioned above, this cannot be obtained in the absence of noise, see [60].
For passive scalars a version of Theorem 1.1 holds in all dimensions.
@@
Let $d\in{\mathbb{N}}_{\geq 1}$, $N\geq 1$ and $\delta>0$ be fixed. Then the following hold.
  1. (1)
@@
    (Small noise). For all $\mu>0$, there exists a family $(\theta^{\gamma})_{\gamma\in(0,1)}\subseteq\mathcal{S}_{\ell^{2}}^{0}$ such that, for all mean-zero $\varrho_{0}\in H^{\delta}(\mathbb{T}^{d})$ satisfying $N^{-1}\leq\|\varrho_{0}\|_{L^{2}}$ and $\|\varrho_{0}\|_{H^{\delta}}\leq N$, we have
    $\inf_{\gamma\in(0,1)}\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{d}}\gamma\,|\nabla\varrho^{\gamma}|^{2}\,\mathrm{d}x\,\mathrm{d}t>0.$
  2. (2)
    (Prescribed rate). For all $\eta\in(0,1)$ and $\mu>-\ln(1-\eta)$, there exists a family $(\theta^{\gamma})_{\gamma\in(0,1)}\subseteq\mathcal{S}_{\ell^{2}}^{0}$ such that, for all mean-zero $\varrho_{0}\in H^{\delta}(\mathbb{T}^{d})$ satisfying $N^{-1}\leq\|\varrho_{0}\|_{L^{2}}$ and $\|\varrho_{0}\|_{H^{\delta}}\leq N$, we have
@@
    $\inf_{\gamma\in(0,1)}\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{d}}\gamma\,|\nabla\varrho^{\gamma}|^{2}\,\mathrm{d}x\,\mathrm{d}t\geq\frac{\eta}{2}\,\|\varrho_{0}\|_{L^{2}}^{2}.$
In the statements in (1)-(2), $\varrho^{\gamma}$ denotes the unique global smooth solution to (1.3).
@@
The comments below Theorem 1.1 extend to the above result, while the existence of a global smooth solution of (1.3) with $\theta^{\gamma}\in\mathcal{S}_{\ell^{2}}^{0}$ follows from, e.g., [6, Theorem 4.2]. In particular, arguing as in (1.5), Theorem 1.2(1) yields anomalous dissipation for passive scalars even in presence of a small transport noise.
@@
Next, we provide a heuristic motivation for the $\nu$-dependence of the transport noise in the context of NSEs with transport noise (1.1) and a possible physical interpretation of our results.

### v3

@@ Abstract.
In this paper, we show that suitable transport noises produce anomalous dissipation of both enstrophy of solutions to 2D Navier-Stokes equations and of energy of solutions to diffusion equations in all dimensions. The key ingredients are Meyers’ type estimates for SPDEs with transport noise, which are combined with recent scaling limits for such SPDEs. The former enables us to establish, for the first time, uniform-in-time convergence in a space of positive smoothness for such scaling limits. Compared to previous work, one of the main novelties is that anomalous dissipation might take place even in the presence of a transport noise of arbitrarily small intensity. Physical interpretations of our results are also discussed.
@@ 1. Introduction and statement of the main results
(1.1) $\left\{\begin{aligned} \partial_{t}\zeta^{\nu}&+(u^{\nu}\cdot\nabla)\zeta^{\nu}=\nu\Delta\zeta^{\nu}+\sqrt{2\mu}\sum_{k\in{\mathbb{Z}}^{2}_{0}}\theta_{k}^{\nu}\,(\sigma_{k}\cdot\nabla)\zeta^{\nu}\circ\dot{W}_{t}^{k}&\text{on }&\mathbb{T}^{2},\\ \zeta^{\nu}(0,\cdot)&=\zeta_{0}&\text{on }&\mathbb{T}^{2},\end{aligned}\right$.
where $u^{\nu}={\mathcal{K}}\,\zeta^{\nu}\stackrel{{\scriptstyle{\rm def}}}{{=}}(-\partial_{y},\partial_{x})\Delta^{-1}\zeta^{\nu}$, $\mathcal{K}$}\, denotes the Biot-Savart operator, $\mathbb{T}^{2}=({\mathbb{R}}/{\mathbb{Z}})^{2}$ the two-dimensional torus, $\nu>0$ the kinematic viscosity of the fluid, $(W^{k})_{k\in{\mathbb{Z}}^{2}_{0}}$ a family of complex Brownian motions, $\circ$ the Stratonovich product, $\theta^{\nu}=(\theta_{k}^{\nu})_{k\in{\mathbb{Z}}^{2}_{0}}\in\ell^{2}$ with ${\mathbb{Z}}^{2}_{0}={\mathbb{Z}}^{2}\setminus\{0\}$, $(\sigma_{k})_{k\in{\mathbb{Z}}^{2}_{0}}$ a family of divergence-free vector field described in Subsection 2.2, and $\mu>0$ the noise’s intensity (cf. (1.5) below).
Anomalous energy dissipation in fluid flows has been proven experimentally to a large degree [85] and stands at the basis of turbulence theory [48, 59, 60, 61]. For this reason, it is sometimes referred to as the zeroth law of turbulence. Roughly speaking, anomalous dissipation states that at high Reynolds numbers (i.e., $\nu\downarrow 0$) the averages of the energy dissipation are uniformly bounded from below in $\nu$. In the context of the Kraichnan-Batchelor theory (KB in the following) of 2D turbulence [15, 62] (see [48, Section 9.7] for additional references), the relevant anomalously dissipated quantity is the enstrophy, and the anomalous dissipation of enstrophy reads as:
@@
In the above, the (unspecified) operator $\langle\cdot\rangle$ typically represents an ensemble average, e.g., space-time average or an expected value of it in the case of random environments. The physical mechanism behind the anomalous dissipation of enstrophy is the transfer, as $\nu\downarrow 0$, of the enstrophy from large to small spatial scales by the nonlinear convective term. This transference produces high gradients $|\nabla\zeta^{\nu}|^{2}$ which cannot be compensated by the small multiplicative factor $\nu$ and therefore yield a non-trivial limit energy dissipation rate as $\nu\downarrow 0$. Let us mention that, sometimes in the literature, the $\liminf_{\nu\downarrow 0}$ in (1.2) is replaced by a $\limsup_{\nu\downarrow 0}$, resulting in a weaker notion that only guarantees anomalous dissipation of solutions to the 2D NSEs (1.1) along some subsequence of viscosities $\nu=\nu^{k}\to 0$.
In the deterministic setting (i.e., $\theta^{\nu}\equiv 0$), there are serious obstructions in making the above pictures rigorous. Indeed, as explained in [66, Section 1], in the case of $\zeta_{0}\in L^{2}(\mathbb{T}^{2})$, the anomalous dissipation of enstrophy (1.2) implies the existence of enstrophy-dissipative (suitable) weak solutions to the deterministic incompressible 2D Euler equations obtained via vanishing viscosity limits. However, the latter is known not to be true, see e.g., [30, 77] or [28, Remark 2]. Moreover, the proposed ways to fix such a paradox in [35] appeared not to be resolutive [66, Section 7].
The anomalous dissipation result for the enstrophy for the 2D NSEs (1.1) in Theorem 1.1 below appears to be the first result in the direction of the KB theory in the physically relevant case of $L^{2}(\mathbb{T}^{2})$ initial enstrophy, see also the comments below (1.6). However, here we are not aiming at capturing the sophisticated mechanics behind anomalous dissipation as described below (1.2). Instead, we prove that transport noises, localised at high frequencies and varying with the viscosity, lead to anomalous dissipation. More precisely, anomalous dissipation is created by the ‘turbulent flows’ $\sum_{k\in{\mathbb{Z}}^{2}_{0}}\theta^{\nu}_{k}\sigma_{k}\,\dot{W}^{k}$ rather than the convective nonlinearity as expected in the KB theory. Moreover, this is also true for noises of ‘small intensities’ and for ‘nonlinear passive scalar’ such as the 2D NSEs in vorticity formulation (here, we mean the turbulent flows differ from the advected quantity). Surprisingly, stochastic perturbations of 2D NSEs provide new dissipation mechanics compared to the deterministic case. The reader is also referred to Section 5, where we also discuss the velocity formulation of the NSEs and the three-dimensional case. Further details about the physical interpretations of our results are given in Subsection 1.2.
@@
(1.3) $\left\{\begin{aligned} \partial_{t}\varrho^{\gamma}&=\gamma\Delta\varrho^{\gamma}+\sqrt{c_{d}\mu}\sum_{k\in{\mathbb{Z}}^{d}_{0}}\sum_{1\leq\alpha\leq d-1}\theta_{k}^{\gamma}\,(\sigma_{k,\alpha}\cdot\nabla)\varrho^{\gamma}\circ\dot{W}_{t}^{k,\alpha}&\text{on }&\mathbb{T}^{d},\\ \varrho^{\gamma}(0,\cdot)&=\varrho_{0}&\text{on }&\mathbb{T}^{d}.\end{aligned}\right$.
@@
Here, $\mathbb{T}^{d}=({\mathbb{R}}/{\mathbb{Z}})^{d}$ is the $d$-dimensional torus, $\gamma>0$ the diffusivity of the scalar, $\theta^{\nu}=(\theta_{k}^{\gamma})_{k\in{\mathbb{Z}}^{d}_{0}}\in\ell^{2}$, $c_{d}=\frac{d}{d-1}$, while $(W^{k,\alpha})_{k,\alpha}$ and $(\sigma_{k,\alpha})_{k,\alpha}$ are families of complex Brownian motions and divergence-free vector fields described in Subsection 2.2.
@@
Anomalous dissipation for passive scalars advected by turbulent flows has attracted a lot of interest in recent years. For passive scalars, anomalous dissipation by turbulent flows is at the basis of the corresponding theory of scalar turbulence [32, 83, 84]. It can be defined by replacing in (1.2) the viscosity $\nu$ and the vorticity $\zeta^{\nu}$ by the diffusivity $\gamma>0$ and the intensity of the passive scalar $\varrho^{\gamma}$ (solving an advection-diffusion such as (1.3)), respectively. In the deterministic setting, many results have been established in the context of anomalous dissipation. It is not possible to give here a complete overview of the deterministic results, and the reader is referred to, e.g., [14, 19, 23, 27, 29, 34, 52] and the references therein.
@@
Before discussing works on anomalous dissipation and related topics in the context of stochastic fluid dynamics, let us first discuss the physical relevance of transport noise in stochastic fluid dynamics. Nowadays, there are several derivations of NSEs with transport noise available in the literature, see e.g., [31, 46, 47, 56, 70, 73], and it is by now a well-established model in stochastic fluid dynamics [20, 21, 22, 26, 40, 45, 53, 54]. To some extent, at the basis of such derivations is the idea of the separation of scales. A heuristic derivation using the latter principle is given in Subsection 1.2 below. This allows us to stress the physical relevance of the $\nu$-dependence of transport noise in (1.1) and provide an interpretation of our results for 2D NSEs. Starting from the works [42, 49], the effect of transport noise on mixing and enhanced dissipation of NSEs or advection-diffusion equations is by now well-understood, see e.g., [39, 38, 67] (and also [1, 24, 36, 41, 43, 44, 64, 68] for related works). The same is not true for anomalous dissipation. To the best of our knowledge, the only results on the anomalous dissipation for SPDEs with transport noise are given in [55, 80]. A comparison is postponed to Subsection 1.3.4, where a more detailed discussion on our contribution is possible. Finally, from a technical point of view, our proofs rely on a refinement of the scaling limit arguments [42, 49] in the parabolic setting, which is interesting on its own. A detailed discussion is given in Subsection 1.3.2 below.
@@ 1.1. Anomalous dissipation results
In this subsection, we state the main result of the manuscript. Below, for convenience, we employ the following notation
@@
Let $N\geq 1$ and $\delta>0$ be fixed. Then, for all $\mu>0$, there exists a family $(\theta^{\nu})_{\nu\in(0,1)}\subseteq\mathcal{S}_{\ell^{2}}^{0}$ such that, for all mean-zero $\zeta_{0}\in H^{\delta}(\mathbb{T}^{2})$ satisfying $\|\zeta_{0}\|_{H^{\delta}(\mathbb{T}^{2})}\leq N$, we have
$\liminf_{\nu\downarrow 0}\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{2}}\nu\,|\nabla\zeta^{\nu}|^{2}\,\mathrm{d}x\,\mathrm{d}t\geq\frac{1}{2}(1-e^{-\mu/(4\pi^{2})})\,\|\zeta_{0}\|_{L^{2}(\mathbb{T}^{2})}^{2},$
where $\zeta^{\nu}$ is the unique global smooth solution to (1.1).
Solutions to (1.1) are defined in Definition 2.2. The existence of global-in-time and smooth-in-space solutions to (1.1) with $\theta^{\nu}\in\mathcal{S}_{\ell^{2}}^{0}$ follows from the arguments presented in [10, Theorems 2.7 and 2.12]. Under additional assumption on $\zeta_{0}$, Theorem 1.1 also holds with $\inf$ in place of $\liminf$; see Proposition 4.6.
@@
Theorem 1.1 shows that, with an appropriate choice of the noise coefficient $\theta^{\nu}$, anomalous dissipation for the 2D NSEs (1.1) can occur even in the presence of small noise (i.e., for any $\mu>0$). This stands in contrast to the existing literature, which typically requires noise of sufficiently large intensity to observe anomalous dissipation (see Subsection 1.3.4). The main novelty of this work lies in showing that anomalous dissipation can arise even when the transport noise has arbitrarily small intensity. Finally, we note that Theorem 1.1 also allows one to prescribe the fraction $\eta\in(0,1)$ of the initial energy $\|\zeta_{0}\|_{L^{2}}^{2}$ that is anomalously dissipated by the 2D NSEs (1.1). Specifically, by choosing $\mu=-\ln(1-\eta)$, Theorem 1.1 guarantees that the energy dissipation $\liminf_{\nu\downarrow 0}\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{2}}\nu\,|\nabla\zeta^{\nu}|^{2}\,\mathrm{d}x\,\mathrm{d}t$ is at least $\frac{\eta}{2}\|\zeta_{0}\|_{L^{2}(\mathbb{T}^{2})}^{2}$.
To connect Theorem 1.1 with the discussion around (1.2) on the KB theory, let us discuss its implications. Firstly, as $\nabla\cdot\sigma_{k,\alpha}=0$, solutions to (1.1) satisfy:
If $\theta^{\nu}$ is chosen as in Theorem 1.1, then the 2D NSEs (1.1) exhibit anomalous dissipation of enstrophy at time $t=1$, i.e., $\liminf_{\nu\downarrow 0}\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{2}}\nu\,|\nabla\zeta^{\nu}|^{2}\,\mathrm{d}x\,\mathrm{d}t>0,$ or equivalently,
Thus, with the aid of transport noise, we obtain anomalous dissipation of enstrophy while the energy balance (1.6) remains true at fixed $\nu>0$ and with $\zeta_{0}\in L^{2}(\mathbb{T}^{2})$. As mentioned above, this cannot be obtained in the absence of noise, see [66].
An inspection of the proof of Theorem 1.1 shows that the time $t=1$ can be replaced by any time $t>0$; however, the corresponding choice of $(\theta^{\nu})_{\nu\in(0,1)}$ depends on such time $t$. In addition, we can take $\theta^{\nu}$ to be constants for $\nu\in[\nu_{j+1},\nu_{j})$ where $(\nu_{j})_{j\geq 1}\subseteq(0,1]$ is a sequence satisfying $\nu_{1}=1$, $\nu_{j+1}\leq\nu_{j}$ and $\lim_{j\to\infty}\nu_{j}=0$. Finally, the proof of Theorem 1.1 shows that $\theta^{\nu}$’s are localized at high frequencies:
@@
For passive scalars, a version of Theorem 1.1 holds in all dimensions.
@@
Let $d\in{\mathbb{N}}_{\geq 1}$, $N\geq 1$ and $\delta>0$ be fixed. Then, for all $\mu>0$, there exists a family $(\theta^{\gamma})_{\gamma\in(0,1)}\subseteq\mathcal{S}_{\ell^{2}}^{0}$ such that, for all mean-zero $\varrho_{0}\in H^{\delta}(\mathbb{T}^{d})$ satisfying $\|\varrho_{0}\|_{H^{\delta}}\leq N$, we have
@@
$\liminf_{\gamma\downarrow 0}\mathbf{E}\int_{0}^{1}\int_{\mathbb{T}^{d}}\gamma\,|\nabla\varrho^{\gamma}|^{2}\,\mathrm{d}x\,\mathrm{d}t\geq\frac{1}{2}(1-e^{-\mu/(4\pi^{2})})\,\|\varrho_{0}\|_{L^{2}(\mathbb{T}^{d})}^{2},$
where $\varrho^{\gamma}$ is the unique global smooth solution to (1.3).
Global solutions to (1.3) are defined in Definition 2.1, and their existence as well as smoothness for $\theta^{\gamma}\in\mathcal{S}_{\ell^{2}}^{0}$ is well-known, e.g., [7, Theorem 4.2]. Under an additional assumption on $\varrho_{0}$, the $\liminf_{\gamma\downarrow 0}$ can replaced by $\inf_{\gamma\in(0,1)}$ in Theorem 1.2, see Proposition 3.6.
@@
The comments below Theorem 1.1 extend to the above result. In particular, arguing in the comments below (1.5), Theorem 1.2 yields anomalous dissipation for passive scalars even in the presence of a small transport noise.
Remark 1.3.
Here we collect some additional comments on the anomalous dissipation of passive scalars as in Theorem 1.2. The same consideration extends verbatim to the 2D NSEs case analyzed in Theorem 1.1.
  • (Limiting behavior of $(\varrho^{\gamma})_{\gamma\in(0,1)}$). From the proof of Theorem 1.2, it follows that the unique global solution $\varrho^{\gamma}$ to (1.3) with $\theta^{\gamma}$ as in the latter satisfies
    (1.8) $\varrho^{\gamma}\to\varrho_{\mathrm{det}}^{0}\ \text{as $\gamma\downarrow 0$ in probability in }C([0,1];L^{2}(\mathbb{T}^{d})),$
    where $\varrho_{\mathrm{det}}^{0}$ is the unique global solution to
    (1.9) $\partial_{t}\varrho_{\mathrm{det}}^{0}=\mu\Delta\varrho_{\mathrm{det}}^{0}\ \text{ on }\mathbb{T}^{d},\qquad\varrho_{\mathrm{det}}^{0}(0,\cdot)=\varrho_{0}\ \text{ on }\mathbb{T}^{d}.$
    This fact is crucial in our proof of Theorem 1.2, see Subsection 1.3.1 for more details. Furthermore, the appearance of the additional (eddy) dissipation $\mu\Delta\varrho_{\mathrm{det}}^{0}$ in (1.9) underlies the anomalous dissipation of energy for passive scalars. It is worth noticing that (1.8) implies that the extra dissipation in (1.9) is produced by the gradients of $\varrho^{\gamma}$. More precisely, (1.8) and the energy balances for $\varrho^{\gamma}$ and $\varrho_{\mathrm{det}}^{0}$ yield
    $2\int_{0}^{\cdot}\int_{\mathbb{T}^{d}}\gamma\,|\nabla\varrho^{\gamma}|^{2}\,\mathrm{d}x\,\mathrm{d}s=\|\varrho_{0}\|_{L^{2}(\mathbb{T}^{d})}^{2}-\|\varrho^{\gamma}(\cdot)\|_{L^{2}(\mathbb{T}^{d})}^{2}$
    $\stackrel{{\scriptstyle\gamma\downarrow 0}}{{\to}}\|\varrho_{0}\|_{L^{2}(\mathbb{T}^{d})}^{2}-\|\varrho_{\mathrm{det}}^{0}(\cdot)\|_{L^{2}(\mathbb{T}^{d})}^{2}$ $=2\int_{0}^{\cdot}\int_{\mathbb{T}^{d}}\mu\,|\nabla\varrho_{\mathrm{det}}^{0}|^{2}\,\mathrm{d}x\,\mathrm{d}s$
    in probability in $C([0,1])$. The reader is referred to Remark 3.4 for an analogous result in the diffusive case $\gamma>0$.
    Let us emphasize that the convergence in (1.8) is non-trivial due to the vanishing diffusive limit $\gamma\downarrow 0$. While compactness and the energy balance yield $\varrho^{\gamma}\to\varrho_{\mathrm{det}}^{0}$ in $C([0,1];H^{-\delta}(\mathbb{T}^{d}))$ for all $\delta>0$, establishing the stronger convergence in (1.8) requires a more delicate argument. In particular, (1.8) relies on a careful choice of the coefficients $\theta^{\gamma}$, which is made possible by a scaling limit argument and the application of Meyers’ estimates (proved in Appendix A). Further details are provided in Subsection 1.3.1.
  • (Non necessity of $L^{2}$-convergence for anomalous dissipation). Although we will prove Theorem 1.2 using the convergence stated in (1.8), we emphasize that this convergence is not required for anomalous dissipation to occur. In fact, it suffices that $\varrho^{\gamma}\approx\varrho_{\mathrm{det}}^{\gamma}$ with high probability for all $\gamma>0$. This observation leads to a slightly different version of the anomalous dissipation result found in Theorems 1.1 and 1.2, which we will present for passive scalars in Proposition 3.6, and for the 2D NSEs in Proposition 4.6.
@@
Next, we provide a heuristic motivation for the $\nu$-dependence of the transport noise in the context of NSEs with transport noise (1.1). We also offer a possible physical interpretation of our results. The physical motivations for the $\gamma$-dependence of the transport noise in advection-diffusion equations (1.3) are not discussed here; we refer the reader to [33, Section 2] and [80, Section 2] for further discussion.